OpenStudy (anonymous):

What is the area of this regular polygon? A. 16.8 ft2 B. 23.4 ft2 C. 46.8 ft2 D. 49.2 ft2

OpenStudy (igreen):

Picture?

OpenStudy (igreen):

That is a hexagon, we can use the formula: \(A = \dfrac{3 \sqrt 3}{2} a^2\) Where 'a' is the side length. So plug in 3 for 'a': \(A = \dfrac{3 \sqrt 3}{2} (3^2)\) Simplify \(3 \sqrt 3\): \(A = \dfrac{5.19615242}{2} (3^2)\) Simplify

OpenStudy (igreen):

@Kamisia

OpenStudy (anonymous):

Yes?

OpenStudy (igreen):

Simplify that

OpenStudy (igreen):

\(3^2 =~?\)

OpenStudy (anonymous):

9

OpenStudy (igreen):

Yes, now divide 5.19615242 / 2

OpenStudy (anonymous):

2.59807621

OpenStudy (igreen):

Yes, now multiply 9 * 2.59807621

OpenStudy (anonymous):

23.38268589

OpenStudy (igreen):

Yes, and that rounds to 23.4.

OpenStudy (igreen):

That's our answer :)

OpenStudy (anonymous):

Thanks @iGreen

OpenStudy (igreen):

Oh, and \(\bbox [10pt, lime, border:5pt solid black]{\huge\cal\color{red}\succ\color{blue}{Welcome\ to\ \color{#00A1FF}{Open}\color{#329932}{Study!\color{red}\prec}}}\)

OpenStudy (igreen):

No problem

OpenStudy (vampireprincess):